OpenStudy (anonymous):

simplify:(cos^4x-sin^4x)/(cos^2x-sin^2x)

OpenStudy (joannablackwelder):

I would factor the numerator by difference of squares rule and cancel.

OpenStudy (anonymous):

ive been trying to change sin^2x into 1+cos^2x

OpenStudy (joannablackwelder):

I think sin^2x=1-cos^2x

OpenStudy (joannablackwelder):

But I think factoring is easier.

OpenStudy (anonymous):

so (cosx)(cosx)-(sinx)(sinX)

OpenStudy (joannablackwelder):

I would factor the numerator (cos^2x)^2-(sin^2x)^2=(cos^2x-sin^2x)(cos^2x+sin^2x)

OpenStudy (anonymous):

I was woring on it and im confused I think the answer is 2 cos^2x-1 but im not sure

OpenStudy (joannablackwelder):

That's not what I get. Do you see what I did in my last comment?

OpenStudy (anonymous):

yes im sry absolutely suck at this im trying

OpenStudy (joannablackwelder):

You'll get better if you keep trying. :)

OpenStudy (joannablackwelder):

Once you factor the numerator you have a common term in the numerator and denominator. Do you see it?

OpenStudy (anonymous):

yess

OpenStudy (joannablackwelder):

Once you cancel, you are left with only cos^2x+sin^2x which is the pythagorean identity.

OpenStudy (joannablackwelder):

So, it simplifies to 1 :)

OpenStudy (anonymous):

thaks I have 33 problems to do by 8 am and I only have 13 done :(

OpenStudy (joannablackwelder):

Oh no! Let me know if you have more questions or if you get stuck.

OpenStudy (joannablackwelder):

I'll do my best to help :)

OpenStudy (anonymous):

thank you